OpenStudy (anonymous):

Find the value of x so that x, 3x+1 and 5x+1 will form a geometric sequence.

OpenStudy (anonymous):

(3x+1)^2=x(5x+1) 9x^2+6x+1=5x^2+x ==> 4x^2+5x+1=0 x=-1,x=-1/4

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Geometric_sequence